A shopper in a supermarket takes a box of sugar from a shelf that is 1.5 m high because he is going to bake some muffins. If the box has a mass of 1.5 kg, what Potential Energy did the sugar have before it was taken from the shelf?
1
Expert's answer
2020-05-18T10:19:21-0400

Potential energy of the object of mass "m" at height "h" above the ground is "E = m g h".

Hence, potential energy of the sugar on a shelf is "E = 1.5 kg \cdot 9.81 \cdot \frac{m}{s^2} \cdot 1.5 m \approx 22.1 J".